OpenStudy (anonymous):

y = –x + 6 and y = x – 8

OpenStudy (anonymous):

x = 7

OpenStudy (stormfire1):

If you want the steps, set both equations equal to each other and solve for x: -x+6 = x - 8 6 = 2x -8 14 = 2x x = 7